Back up and running

On April 3, 2017, Posted by , In Operations, With Comments Off on Back up and running

We are back up from the long ALS shutdown, and the beamlines are in good shape. We completed a number of projects on the sector 5 beamlines, in particular, as well as performed maintenance on a number of systems:

– Pilatus3 6M installed on 5.0.1. The install went well, and tests with x-rays show that it is working well, except for a column of bad pixels in the lower right-hand corner. We have masked out the bad pixels so they will not affect data processing, and plan to have Dectris come out during the June 2-bunch period to replace the bad module.
– A beta version of B4 has been tested and now available for users and we think intuitive to use, though of course contact staff for a tour of the new GUI anytime. Mounting, Quick Centering and Collection tabs are consolidated into one main panel. Users can also expand additional functionality using the + button. There are a few features in B4 that are not in B3. For instance, the automated data processing results are available through the B4 GUI under the new Data tab.
– A pin base warmer is now installed on 5.0.1 and 5.0.2. For initial implementation, we have the warmer automatically turn on during an exchange, warming the gonio magnet while the pin is not on the magnet. Temperature measurements and diffracting tests show that warming the base itself for two minutes does not result in heating at the loop, and manual warming of the pin base is available through the GUI, for those who would like to use it.
– A new raid system was installed to automatically back up the user home directories (processed data). Note that diffraction images are still automatically deleted after two weeks, unless you contact us.
– A new dewar lid for 5.0.2 (to reduce ice build up) was built and will be installed in our effort to reduce ice buildup.
– The 5.0.2 ion pump was replaced. Even after, we were still getting intermittent trips due to pressure bursts. We reopened and closed the mono tank again thinking that o-ring may be the source of the problem. Since then we haven’t had a trip due to pressure bursts.
– Automatic COR detection scripts were installed, along with a dedicated position in the robot dewar for the alignment pin. Note that currently this is a staff-only feature, but we plan to allow users access to the feature after more testing.
– We are also thoroughly testing an automatic centering of the loop.
– We received from Mitegen and are now installing the new DBS on all of our beamlines. Continuous readout of the beam intensity measure by the DBS during data collection is now available.
– We are implementing a new script to identify the amount of beam that would not saturate the vortex detector. The vortex detector is available for users to do a scan and/or to identify elements in a crystal.

Maintenance activities:

– All cryojets were serviced

– The chillers for 501/503 were flushed

– All detector chillers coolants were replaced

– Motors for dewar motion were replaced

– The viewing prism on 503 was replaced

Thank you for reading!

Comments are closed.